-2x-8y=-42
-4x+6y=-18
1
Expert's answer
2011-05-12T05:58:11-0400
-2x& - 8y = -42
-4x + 6y = -18

-4x& - 16y = -84
-4x& + 6y& = -18

22y = 66
y= 3

-4x& + 6*3& = -18
-4x = -18 - 18 = - 36
x = 9

Answer:& (9,3)